# IGDrivSim: A Benchmark for the Imitation Gap in Autonomous Driving

### v1

@@ III-E Reward
To compute the RL component of the combined loss defined in Equation 1, we utilize the reward function defined in the Waymax simulator, which is the sum of safety penalties incurred during each time step, specifically due to off-road driving and vehicle overlap:
$r_{t}=-\mathds{1}_{\text{overlap}}(t)-\mathds{1}_{\text{off-road}}(t),$
where $\mathds{1}_{\text{overlap}}(t)$ is an indicator function that equals one if the ego vehicle collides with another vehicle at timestep $t$, and $\mathds{1}_{\text{off-road}}(t)$ equals one if the ego vehicle is driving off the road at timestep $t$.
IV Experiments
Our experiments reveal BC’s limitations when used to train driving policies from human demonstrations, and establish IGDrivSim as a valuable benchmark for addressing imitation gaps between human drivers and self-driving agents.
IV-A Training
@@
When training agents with BC, we use two different encoders to embed the scene observations: an MLP for encoding the objects’ features and a polyline encoder, based on the architecture proposed in [23], for the roadgraph. We apply layer norms on the two embeddings, which are then concatenated and passed to a simple recurrent neural network (RNN) with a final softmax layer to compute action probabilities. Supported by the results of [15], we discretize the action space into 256 bins. The RNN enables handling the partial observability in the environment by maintaining and updating a hidden state that captures information about the environment over time despite limited observations at each timestep. For RL, we used the same architecture but added an MLP value head after the RNN cell.
@@
During training with the combined BC and RL loss, we consider two types of trajectories for each scenario: a logged trajectory, where the simulator is updated using the expert’s actions, and a simulated trajectory, where the simulator is updated using the imitator’s actions. The BC component is calculated by computing the negative log-likelihood of the expert’s actions based on the action distribution predicted by the imitator policy over the entire logged trajectory. Furthermore, to reduce distribution shift, we follow a similar approach to [24] by discretizing the expert’s demonstrations using the expert agent introduced in Waymax. The scenario is then reset, and the imitator policy is used to generate a new trajectory, with rewards recorded at each timestep. The RL component, represented by the PPO loss, is computed on this imitator-driven trajectory. Finally, the model is updated using the Adam optimizer with a cosine decay learning rate, ranging from $10^{-3}$ to 00, applied to the combined loss as defined in Equation 1. Each training run took approximately 2 days on a cluster of 8 NVIDIA GTX 1080 GPUs.
@@
IV-B Results
To demonstrate that under imitation gaps—widened in IGDrivSim by artificial constraints on the imitator’s observability—BC alone fails to learn efficient and safe behavior for the imitator’s specific observability, we train imitators with varying observability. We then compare the policies learned through BC on human expert demonstrations with those trained using a combined BC and RL loss. As our focus is on IL of a single agent; while Waymax allows for controlling all the agents of the scene, we restrict the control to the ego car from which the scene was recorded.
@@ IV-B Results
We evaluate four imitators: one with full observability using the initial Waymax settings, two with circular fields of view in IGDrivSim with radii of 4 meters and 6 meters, and one operating under Gaussian noise in IGDrivSim with a standard deviation of 3 meters, as defined in section III-C. These agents’ perceptions differ significantly from human expert perception, thereby introducing an imitation gap in the training process. Notably, even the agent with full observability relies solely on sensor-recorded data, which inherently diverges from true human perception that could include additional factors such as traffic sounds. Our preliminary experiments show that for the observability constraints to impact performance compared to full observability, the radius must be narrow enough to limit the agent’s view of its immediate surroundings and the noise to be strong enough to introduce uncertainty. These restrictions prevent the agent from anticipating changes in the environment and forces it to learn a different policy compared to when it has access to the entire Waymax state, thereby introducing varying degrees of imitation gap.
@@
Table I provides detailed values for the trajectory and per-step metrics—log divergence, overlap, and off-road—while Figure 3 shows a comparison of the per-step metrics, for policies trained with BC and the combined BC-RL loss across the four considered imitators. For all the imitators, balancing the BC loss with the RL loss using the simple penalty reward helps the agent learn a safer policy, reducing both overlap and off-road incidents. This demonstrates that using the BC-RL combined loss allows for better behavior learning under the imitator’s POMDP than BC alone. Although [16] highlights similar findings, they do not address the imitation gap, which we consider a key factor in explaining these results. Moreover, unlike [16], we have open-sourced our code, making these findings more accessible to the research community.
Interestingly, among the imitators with full observability, Gaussian noise, and a circular field of view with a radius of 6, the policy learned with the combined loss not only minimizes safety metrics but also exhibits better log divergence—meaning it is closer to the expert’s policy—than the policy trained only with BC. By contrast, the imitator with a circular field of view of radius 4, also trained with the combined loss, shows higher log divergence than the BC policy, despite improvements in overlap and off-road incidents. This indicates that the policy trained with the combined loss diverges from the expert’s but performs better within its own POMDP. By learning behaviors not demonstrated by the human expert, it effectively reduces overlap and off-road incidents based on its specific perception of the cars and road environment. This result highlights how RL regularization can allow learning optimal behaviors not shown in expert demonstrations, thus mitigating the imitation gap.
Figure 3: Comparison of per-step metrics—log divergence, off-road, and overlap—between policies trained with BC (blue) and the combined BC-RL (BC+RL, orange) using parameters ($w_{BC}=1,w_{RL}=0.05$). The metrics are averaged over two seeds for four imitators defined by different partial observability in IGDrivSim.
TABLE I: Mean trajectory and per-step metrics for policies trained with BC and the combined BC-RL loss under different partial observability in IGDrivSim. Blue cells represent BC-only policies, while orange cells indicate BC-RL policies ($w_{BC}=1,w_{RL}=0.05$). All values are averaged over two seeds.
Metric Full obs Gaussian Circular Circular
noise ($\sigma$=3) FoV (r=4) FoV (r=6)
log 3.67 3.69 4.20 3.87
divergence 3.78 3.85 4.12 3.95
max log 10.56 10.72 12.64 11.51
divergence 11.00 11.25 12.28 11.76
overlap 4.78 4.66 5.32 5.14
(%) 5.02 5.06 5.51 5.33
overlap 21.18 20.91 24.45 22.97
rate (%) 22.20 22.39 24.42 23.61
offroad 2.20 2.36 0.72 1.26
(%) 2.69 3.04 2.18 1.74
offroad 6.87 7.20 4.16 5.01
rate (%) 8.18 8.95 7.31 6.03
@@
Qualitatively, the agent trained with the BC-RL combined loss shows greater reactivity to changes in its receptive field. For instance, it engages in heavy braking when the road lines of other vehicles enter its view. In contrast, in several manually inspected scenarios, expert human drivers, having broader perception, anticipate road limits and nearby vehicles earlier, leading to smoother, more controlled braking and overall fluid driving behavior. The RL regularization of the BC loss enables the agent to learn behaviors that are effective within the imitator’s POMDP but not observed in the expert demonstrations. As a result, as shown in Figure 4, the self-driving car follows trajectories that diverge significantly from the expert-logged paths while remaining safe and efficient.
Figure 4: Three examples showing the imitator (with a circular receptive field of radius 4) trained with the combined BC-RL loss, following distinct paths (in blue) compared to the human expert’s trajectories (in cyan).
V Related Work
A variety of methods have been proposed recently to solve the imitation gap, but a lack of standardized benchmarks makes comparison among the algorithms harder. We provide a principled imitation gap benchmark with real-world motivation from learning policies for self-driving cars. A popular category of algorithms for solving the imitation gap combines BC and RL objectives using different weighting schemes. Weihs et al. [12] propose an algorithm that dynamically weights the objectives during training, while Nguyen et al. [25] start from an offline RL algorithm accelerated by the BC objective. Another solution idea is to use DAgger-style [26] online supervision to accelerate learning in an imitation gap setting [13]. Perhaps most closely related to ours is “Learning by Cheating” by Chen et al. [27], where they consider a vision-based driving setting with an imitation gap problem based on the CARLA driving simulator [8]. Vuorio et al. [14] show that prior information about the desired behaviors is always required to solve the imitation gap and propose a Bayesian solution to the problem. The failure of BC in problems with an imitation gap can be seen as a form of causal confounding [28]. When the imitator can solve the problem without deviating from the expert behavior, Ortega et al. [28] and Swamy et al. [29] show that DAgger-style methods can work well in this setting, while Vuorio et al. [30] use a separate inference model.
Learning driving policies is a popular field of research with multiple datasets and benchmarks targeting it. Waymo Open Dataset (WOMD) [10] contains many driving scenarios with log trajectories of human drivers. The Waymax [15] simulator we build upon runs on the scenarios from the WOMD dataset. Argoverse2 [31] provides an alternative to WOMD, while NuPlan [32] is one for Waymax. We chose Waymax due to its fast accelerator-compatible implementation and the large quantity of high-quality data included in WOMD. A closely related benchmark to ours is [33], which builds on WOMD and features partial observability. However, the effect of partial observability in Nocturne is unknown as it is not ablated.

### v2

@@ IV-E Reward
To compute the RL component of the combined loss defined in Equation 1, we utilize the reward function defined in the Waymax simulator, which is the sum of safety penalties incurred during each time step, specifically due to off-road driving and vehicle overlap: $r_{t}=-\mathds{1}_{\text{overlap}}(t)-\mathds{1}_{\text{off-road}}(t),$ where $\mathds{1}_{\text{overlap}}(t)$ is an indicator function that equals one if the ego vehicle collides with another vehicle at timestep $t$, and $\mathds{1}_{\text{off-road}}(t)$ equals one if the ego vehicle is driving off the road at timestep $t$.
V Experiments
Our experiments highlight BC’s limitations in training driving policies in settings with imitation gaps, compared to a method that combines BC and RL. Each experiment in Section V-B ran on an 8-GPU (NVIDIA GTX 1080) cluster for $\approx$ 2 days, using 230 GB RAM.
V-A Training
@@
For BC training, we use an MLP to encode the object features and a polyline encoder from [34] for the roadgraph. After applying layer norms, the embeddings are concatenated and passed to an RNN with a final softmax to predict action probabilities. Following [15], the action space is discretized into 256 bins. The RNN enables handling the partial observability in the environment by maintaining and updating a hidden state that captures information about the environment over time despite limited observations at each timestep. For RL, we used the same architecture but added an MLP value head after the RNN cell.
TABLE I: Regularizing BC with RL improves safety metrics. Mean trajectory and per-step metrics for BC and BC-RL policies under varying partial observability in IGDrivSim. Blue cells denote BC-only policies, orange cells BC-RL ($w_{BC}=1,w_{RL}=0.05$). The mean is averaged over three seeds and significant differences are in bold.
Metric Full obs Gaussian Circular Circular Conic FoV Random
noise ($\sigma$=3) FoV (r=4) FoV (r=6) (r=4, $\alpha=\frac{2}{3}\pi$) masking (p=0.7)
log 3.67 3.69 4.18 3.87 4.60 3.68
divergence 3.79 3.85 4.12 3.94 4.19 3.81
max log 10.56 10.72 12.54 11.51 14.08 10.68
divergence 11.00 11.23 12.28 11.74 12.39 11.16
overlap 4.80 4.66 5.28 5.09 3.70 4.95
(%) 5.06 5.02 5.45 5.31 5.73 5.30
overlap 21.17 20.94 24.25 22.73 24.65 21.70
rate (%) 22.24 22.26 24.30 23.47 25.16 23.00
offroad 2.13 2.39 0.77 1.13 1.10 2.20
(%) 2.54 2.96 2.15 1.68 2.39 2.71
offroad 6.69 7.25 4.28 4.49 7.00 6.85
rate (%) 7.78 8.73 7.18 5.73 8.34 7.96
Figure 3: Comparison of per-step metrics—log divergence, off-road, and overlap—between policies trained with BC (blue) and the combined BC-RL (orange) using parameters ($w_{BC}=1,w_{RL}=0.05$). The metrics are averaged over three seeds, with 95% confidence intervals shown as error bars, for imitators under different partial observability.
@@
During training with the BC-RL combined loss, each scenario includes a logged trajectory—where the environment is updated by expert actions—and a simulated trajectory—where it is updated by imitator actions. The BC loss component is computed on the logged trajectory using the equation defined in Section III-A. To reduce distribution shift, similarly to [26], we discretize the expert’s demonstrations using the expert agent introduced in Waymax. The RL loss component (PPO loss), is computed on the simulated trajectory where reward has been recorded at each timestep. Finally, the model is updated using the Adam optimizer with a cosine decay learning rate, ranging from $10^{-3}$ to 00, applied to the combined loss defined in Equation 1.
@@
V-B Results
To show that BC alone fails to learn efficient and safe behavior under imitation gaps—widened in IGDrivSim by observability constraints—we train imitators with varying observability levels. We then compare the policies learned through BC on human expert demonstrations with those trained using a combined BC and RL loss. As our focus is on IL of a single agent; while Waymax allows for controlling all the agents of the scene, we restrict the control to the ego car from which the scene was recorded.
@@ V-B Results
We evaluate the imitator under full scene observability and five different limited observability constraints: (1) limited field of view—either circular regions with radii of 6 and 4 meters, or a conical region with an angle of $\theta=\frac{2}{3}\pi$ and a radius of 4 meters; and (2) noisy perception—where Gaussian noise (standard deviation: 3 meters) affects the imitator’s position, or the vehicle’s position is randomly masked with probability $p=0.7$ (section IV-C). These agents’ perceptions differ significantly from human expert perception, thereby introducing an imitation gap in the training process. Notably, even the agent with full observability relies solely on sensor-recorded data, which inherently differs from true human perception that could include additional factors like traffic sounds.
@@
Figure 3 compares per-step metrics (log divergence, overlap, and off-road) for policies trained with BC and BC-RL across the six imitators, while Table I provides the detailed values along with the trajectory metrics. For all the imitators, balancing the BC loss with the RL loss using the simple penalty reward helps the agent learn a significantly safer policy, reducing both overlap and off-road incidents. This demonstrates that using the BC-RL combined loss allows for better behavior learning under the imitator’s POMDP than BC alone. Although [16] highlights similar findings, they do not address the imitation gap, which we consider a key factor in explaining these results. Moreover, we have open-sourced our code, making these findings more accessible to the research community.
Furthermore, as shown in Figure 3, the policies trained with the combined BC-RL loss not only improve safety metrics but also achieve lower log divergence—indicating closer alignment with the expert’s policy—compared to those trained solely with BC. The only exception is the imitator with a conic field of view ($\theta=\frac{2}{3}\pi$, radius = 4 meters), which has the strictest observability constraint in our experiment. This indicates that under this partial observability, the policy trained with the combined loss diverges from the expert’s but performs better (i.e., more safely) within its own POMDP. RL regularization allows the imitator to learn behaviors not demonstrated by the human expert but effective in reducing overlap and off-road incidents based on its perception of the cars and road environment, demonstrating how RL helps narrow the imitation gap.
@@
Qualitatively, the agent trained with the BC-RL combined loss shows greater reactivity to changes in its receptive field. For example, it engages in heavy braking when the road lines of other vehicles enter its view. In contrast, in several manually inspected scenarios, expert human drivers, having broader perception, anticipate road limits and nearby vehicles earlier, leading to smoother braking and overall fluid driving behavior. The RL regularization of the BC loss enables the agent to learn effective behaviors within the imitator’s POMDP that were not observed in the expert demonstrations.